How do I work out 2(3x-4)=22
Found 2 solutions by Boreal, macston:
Answer by Boreal(15235)

2(3x-4)=22
distribute
6x-8=22
add 8 to both sides
6x=30
divide by 6 both sides
x=5
===============
check
2(11)=22

Answer by macston(5194)

.
2(3x-4)=22 . Multiply (3x-4) by 2
6x-8=22 . Add 8 to each side.
6x-8(+8)=22(+8)
6x=30 . Divide each side by 6
6x/6=30/6
(6/6)x=5 since 6/6=1:
x=5
.
CHECK: (Plug the answer into the original equation)
2(3x-4)=22
2(3(5)-4)=22
2(15-4)=22
2(11)=22
22=22 This is true, so the solution is correct.